OpenStudy (anonymous):

I don't understand Rational Exponents... could someone explain them to me so I can finish this homework?

OpenStudy (anonymous):

What do you need to know? A rational exponent is just something to the form of \(n^{a/b}\) which is equivalent to \(\sqrt[b]{n^a}\).

OpenStudy (anonymous):

well... i have to divide them and multiply them.... and i'm confused.

OpenStudy (anonymous):

There are some basic rules for division and multiplication of exponents. \[\Large n^a \times n^b = n^{a+b} \\\Large \frac{ n^a}{n^b }=n^{b-a}\]

OpenStudy (anonymous):

Sorry, that should be: \[\Large \frac{n^a}{n^b} = n^{a-b}\]

OpenStudy (anonymous):

so... y2/5 x y^3/8= y^27/40?

OpenStudy (anonymous):

Not quite. 2/5 + 3/8 = 16/40 + 15/40 = 31/40. But I think you're getting it.

OpenStudy (anonymous):

so then y^5/8 / y^1/2 = y^1/8?

OpenStudy (anonymous):

That's right.

OpenStudy (anonymous):

ok, thank you i think i have it ^^

OpenStudy (anonymous):

No problem =)